OpenStudy (anonymous):

Determine all the critical points for the function f(x)= x^2 + (16/x)

OpenStudy (freckles):

You found f' yet?

OpenStudy (anonymous):

f '= 2x-(16/x^2)???

OpenStudy (freckles):

Yep! :)

OpenStudy (freckles):

Now combine fractions! :)

OpenStudy (freckles):

\[f'=\frac{2x(x^2)-16}{x^2}\] right? :)

OpenStudy (anonymous):

Thats what I got! :)

OpenStudy (freckles):

Now set Mr.top=0! :) and solve for x We would look at where the bottom is 0 except our function is not even define for x=0 So we are only look for when the top of f' is 0

OpenStudy (freckles):

\[2x(x^2)-16=0\]

OpenStudy (freckles):

Can you solve this for x?

OpenStudy (anonymous):

x=2 ??

OpenStudy (freckles):

\[2x^3-16=0\] Yes that is the only real critical number \[x^3-8=0 \] \[(x-2)(x^2+2x+4)=0\] x=2 --- \[b^2-4ac=2^2-4(1)(4)=4-16=-12 <0 => x=2 \text{ is the only critical number}\] :) So good job! :)

OpenStudy (anonymous):

Oh great! thank you

OpenStudy (freckles):

You deserve a medal! Seriously great job! :)